OpenStudy (anonymous):

solve the equation for 2^x-5=3

OpenStudy (anonymous):

X=4 2x^2-5=3. Plus 5 on both sides then you get 2^2=8. Then divide by 2 on both sides and your answer will be x=4

OpenStudy (anonymous):

*2x^2=8